OpenStudy (anonymous):

Solve. Find exact solution: (3/x+5)-(11/3x-1)=-1

OpenStudy (anonymous):

is it \[\frac{3}{x+5}-\frac{11}{3x-1}=-1\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

these are a pain in the neck multiply both sides by \[(x+5)(3x-1)\] to get \[3(3x-1)-11(x+5)=-(x+5)(3x-1)\] as a first step

OpenStudy (anonymous):

then multiply out and get \[9x-3-11x-55=-3 x^2-14 x+5\] and go from there

OpenStudy (anonymous):

ok :) whats next

OpenStudy (anonymous):

combine like terms set it equal to zero, solve the resulting quadratic equation

OpenStudy (anonymous):

if you do that correctly, you should get \[3 x^2+12 x-63 = 0\]

OpenStudy (anonymous):

divide by 3 and get \[x^2+4x-21=0\] which fortunately for your factors as \[(x+7)(x-3)=0\]